# Auf der Morgenstelle 14, D-72076 Tübingen, Germany

### v1

@@ 5 Conclusion and Outlook
We have demonstrated that the proposed $x$-space method implemented in POMPOM is capable of accurately solving the DGLAP equation for the non-singlet and singlet cases. We obtained good agreement with both the Les Houches benchmark set as well as the PDFs from Gluck:1998xa . Sum rule violations drop to levels below $10^{-6}$ for $\mathcal{O}(200)$ basis functions, relative corrections when further increasing the number of basis functions are below $10^{-4}$ except very close to $x=1$. POMPOM is accurate over a wide range of $x$ as well as $\mu$, we tested for $2\,\Lambda_{\text{QCD}}\simeq\mu<10\,\mathrm{TeV}$ and $110^{-7}<x<$ 1. The largest deviations are observed for $x\approx 1$ and originate from truncating of the series for $\ln(1-x)$. One could try to improve accuracy in the large-$x$ region by amending the basis by $\ln(1-x)$, however since we achieved a, for our purposes, satisfactory level of accuracy, this is beyond the scope of the present paper.
@@
where $\mathfrak{S}^{\pm}$ and $\mathfrak{F}^{\pm}$ are functions of two momentum fractions. The $\mathbb{H}^{\pm}$ are linear one-dimensional integral operators. Expressions for them can be found in the appendix of Braun2009 . The non-singlet case decouples similar to DGLAP evolution. The existing evolution code t3evol Pirnay:2013 for eq. (43) uses discretization in both $\mu$ and $x$-space. This code was for example applied in Hatta:2019 .
@@ Acknowledgements.
We thank Dominik Bammert for inspiring this work. We are grateful to Werner Vogelsang for useful comments regarding this work. This work has been supported by Deutsche Forschungsgemeinschaft (DFG) through the Research Unit FOR 2926 (project 409651613). J.H. is grateful to the Landesgraduiertenförderung Baden-Württemberg for supporting her research.
@@ Appendix A Master integrals
$I_{1}^{m,n}\,$ $\equiv\,x^{n}\int_{x}^{1}\mathrm{d}\xi\,\frac{\xi^{-n}-1}{1-\xi}\frac{\ln^{m}(x/\xi)}{m!}$ (44)
=$=\,\left\{\begin{array}[]{ll}\sum_{k=0}^{-n-1}\left[\sum_{j=0}^{m}\frac{1}{(k+1)^{m+1-j}}\,\frac{\ln^{j}(x)\,x^{n}}{j!}\,-\,\frac{x^{k+n+1}}{(k+1)^{m+1}}\right]&\quad\text{if }n\,\leq\,-2\,,\\ -\sum_{k=0}^{m}\frac{\ln^{k}(x)}{k!}\,+\,x&\quad\text{if }n\,=\,-1\,,\\ 0&\quad\text{if }n\,=\,0\,,\\ (-1)^{m}\sum_{k=1}^{n-1}\frac{x^{-k}}{k^{m+1}}\,-\,\sum_{k=0}^{m}(-1)^{k}\,H_{n-1,k+1}\,\frac{\ln^{m-k}(x)}{(m-k)!}\,-\,\frac{\ln^{m+1}(x)}{(m+1)!}&\quad\text{if }n\,\geq\,1\,.\end{array}\right.$ (49)
@@
$=\,\sum_{k=0}^{m-1}\zeta_{m-k+1}\frac{\ln^{k}(x)}{k!}\,+\,\frac{\ln^{m}(x)}{m!}\sum_{k=0}^{\infty}\frac{x^{k}}{k}\,-\,\sum_{k=1}^{\infty}\frac{x^{k}}{k^{m+1}}.$ (51)
@@
Here, $H_{n,m}\equiv\sum_{k=1}^{n}\frac{1}{k^{m}}$ are the generalized harmonic numbers, while $\zeta_{m}\equiv\sum_{k=0}^{\infty}\frac{1}{k^{m}}$ are the integer values of the Riemann zeta function.
@@ Appendix C Input distributions and parameters
$\alpha_{s}\!\left(\mu^{2}\right)=\,\frac{4\pi}{\beta_{0}}\,\ln\!\left(\frac{\mu^{2}}{\Lambda_{\text{QCD}}^{2}}\right),$ (73)

### v5

@@ 5 Conclusion and Outlook
We have demonstrated that the proposed $x$-space method implemented in POMPOM is capable of accurately solving the DGLAP equation for the non-singlet and singlet cases. We obtained good agreement with both the Les Houches benchmark set as well as with Mellin-space evolution of the PDFs from Gluck:1998xa . Sum rule violations drop to levels below $10^{-6}$ for $\mathcal{O}(200)$ basis functions, relative corrections when further increasing the number of basis functions are below $10^{-4}$ except very close to $x=1$. POMPOM is accurate over a wide range of $x$ as well as $\mu$, we tested for $2\,\Lambda_{\text{QCD}}\simeq\mu<10\,\mathrm{TeV}$ and $110^{-7}<x<$ 1. The largest deviations are observed for $x\approx 1$ and originate from truncating of the series for $\ln(1-x)$. One could try to improve accuracy in the large-$x$ region by amending the basis by $\ln(1-x)$, however since we achieved a, for our purposes, satisfactory level of accuracy, this is beyond the scope of the present paper. The performed numerical studies demonstrate that for the leading order DGLAP equation the cut-off effect of the POMPOM method can indeed be controlled with a numerically manageable number of basis functions. This is an essential foundation for a potential development of POMPOM towards an state-of-the-art DGLAP evolution program that could for example be used for PDF fitting.
@@
where $\mathfrak{S}^{\pm}$ and $\mathfrak{F}^{\pm}$ are functions of two momentum fractions. The $\mathbb{H}^{\pm}$ are linear one-dimensional integral operators. Expressions for them can be found in the appendix of Braun2009 . The non-singlet case decouples similar to DGLAP evolution. The existing evolution code t3evol Pirnay:2013 for eq. (43) uses discretization in both $\mu$ and $x$-space. [^1] After the publication of the first preprint of the present paper the new twist-3 evolution codes Honeycomb/Snowflake appeared Rodini:2024 . These use an improved discretization approach. This code was for example applied in Hatta:2019 .
@@ Acknowledgements.
We thank Dominik Bammert for inspiring this work. We are grateful to Werner Vogelsang for useful comments during all stages of this work. We also thank Andrea Simonelli for useful discussions regarding the application of the Magnus expansion. The authors acknowledge helpful comments by Johannes Blümlein regarding the literature of state-of-the-art calculations of QCD splitting functions and the running coupling. This work has been supported by Deutsche Forschungsgemeinschaft (DFG) through the Research Unit FOR 2926 (project 409651613). J.H. is grateful to the Landesgraduiertenförderung Baden-Württemberg for supporting her research.
@@ Appendix A Master integrals
$I_{1}^{m,n}\,$ $\equiv\,x^{n}\int_{x}^{1}\mathrm{d}\xi\,\frac{\left(\xi^{-n}-1\right)}{1-\xi}\frac{\ln^{m}(x/\xi)}{m!}$ (44)
=$=\,\left\{\begin{array}[]{ll}-\sum_{k=0}^{-n-1}\left[\sum_{j=0}^{m}\frac{1}{(k+1)^{m+1-j}}\;\frac{\ln^{j}(x)\,x^{n}}{j!}\,-\,\frac{x^{k+n+1}}{(k+1)^{m+1}}\right]&\;\text{if }n\,\leq\,-2\,,\\ -\frac{1}{x}\sum_{k=0}^{m}\frac{\ln^{k}(x)}{k!}\,+\,1&\;\text{if }n\,=\,-1\,,\\ 0&\;\text{if }n\,=\,0\,,\\ x^{n}\left[(-1)^{m}\sum_{k=1}^{n-1}\frac{x^{-k}}{k^{m+1}}\,-\,\sum_{k=0}^{m}(-1)^{k}\,H_{n-1,k+1}\,\frac{\ln^{m-k}(x)}{(m-k)!}\,-\,\frac{\ln^{m+1}(x)}{(m+1)!}\right]&\;\text{if }n\,\geq\,1\,.\end{array}\right.$ (49)
@@
$=\,x^{n}\left[\sum_{k=0}^{m-1}\zeta_{m-k+1}\frac{\ln^{k}(x)}{k!}\,+\,\frac{\ln^{m}(x)}{m!}\sum_{k=1}^{\infty}\frac{x^{k}}{k}\,-\,\sum_{k=1}^{\infty}\frac{x^{k}}{k^{m+1}}\right].$ (51)
@@
Here, $H_{n,m}\equiv\sum_{k=1}^{n}\frac{1}{k^{m}}$ are the generalized harmonic numbers, while $\zeta_{m}\equiv\sum_{k=1}^{\infty}\frac{1}{k^{m}}$ are the integer values of the Riemann zeta function.
@@ Appendix C Input distributions and parameters
$\alpha_{s}\!\left(\mu^{2}\right)=\,\frac{4\pi}{\beta_{0}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\,\ln\!\left(\frac{\mu^{2}}{\Lambda_{\text{QCD}}^{2}}\right)}}\,,$ (73)
@@
Appendix D Computation time
Figure 11: Time required for initialization and evolution of non-singlet and singlet evolution. The timing was performed with Mathematica 13.1 on a ThinkPad T495 with $8\,\mathrm{GB}$ of RAM.
Since the current version of POMPOM is mainly designed for explorative purposes, time performance is not considered the most important aspect of the code. For a potential future upgrade of POMPOM into a code applicable to PDF fits this will be of relevance. One also has to keep in mind that a comparison between evolution programs is only meaningful in reference to a specified task as was dicussed in Candido2022 , where the timing of EKO is compared with APFEL.
To get a first estimate on the time required for PDF evolution with POMPOM, we show initialization times and the time for the evolution to a single scale for non-singlet and singlet evolution in figure 11. Initialization has to be performed only once for a fixed number of basis functions with a certain cut-off to calculate the corresponding evolution matrix. The initialization time grows approximately quadratically with the size of the basis. Due to reuse of already calculated matrix elements for lower number of basis functions, the initialization with higher number of basis functions shown in the upper two plots is slightly quicker than it would be in a standalone calculation. The subsequent evolution to a single scale is much faster in comparison and only grows linearly with increasing number of basis functions. For a reasonable number of basis functions, the evolution to a specific scale takes on the order of $1\,\mathrm{s}$ on a standard consumer laptop.